# Active Gaze Behavior Boosts Self-Supervised Object Learning

### v1

@@ 4 Results
This section first investigates whether toddlers’ gaze behavior during a play session supports learning view-invariant object recognition. We then analyze the factors contributing to the performance of the toddlers’ models.
4.1 Toddler’s central visual experience is curated to develop robust object representations
Here, we evaluate if a toddler’s gaze behavior supports strong object representations. In Figure 4, we compare the representations learned by SimCLR-TT when trained on the different datasets introduced in Section 3.1. We find that models trained with the Toddler fixation dataset outperform those trained with the Random fixation dataset (toddler) or the Centroid fixation dataset (toddler). We also observe that the gap between the Toddler fixation dataset and other datasets built with the ground truth about objects’ identities is small. We see a similar trend with the adult datasets. This suggests that a biologically inspired visual learning model like SimCLR-TT can leverage human gaze behaviors to build better visual representations.
It raises the question of whether the visual experience of only one human during a play session suffices to build good visual representations. To investigate this question, we train one SimCLR-TT on the recordings of each toddler and adult separately and compute the average of linear accuracies. We trained the encoder (ResNet18) using all fixation data from a single toddler/adult, followed by training and testing the linear classifier with the Objects fixation data from the same and different toddlers/adults. We controlled the training set to comprise 75% of the total data, ensuring that the test set did not overlap with the training set. Figure 4 exhibits that the central visual experience of one toddler leads to representations almost as good as those from the central visual experiences of all toddlers.
Finally, we assess whether toddlers’ visual experience produces better or worse representations than that of adults. By comparing the object recognition accuracy of models trained on fixation datasets from toddlers and adults, we see mixed results. Toddlers’ experiences induce more robust representations compared to adults when training with one toddler/adult trial, but not when training with the entire dataset. We conclude that, toddlers’ central visual experience supports more data-efficient learning than adults. Overall, toddlers appear to curate their gaze behavior to develop robust object representations quickly.
Figure 4: Linear object recognition accuracy of training on individual participants and all participants. For gaze fixation, we compared the test results of the model trained on each participant and across all participants. Specifically, the mean accuracy is shown, when trained with 38 dyads individually. For the all-participants setting, 30 participants are randomly selected from the pool of 38 for the training set under each random seed. The two gray bars indicate the test results on the Plain background and Objects fixation datasets. The vertical bars represent the standard deviation over three random seeds.
@@ 4.1 Toddler’s central visual experience is curated to develop robust object representations
4.2 Constrained size of the central visual field is crucial for learning object representations
Previous computational studies neglect the importance of the constrained size of the central visual field for learning visual representations (Orhan et al., 2020; Sheybani et al., 2024). Here, we assess whether our simulated central visual experience leads to better/worse object representations than a wide field of view. We vary the crop size applied to the datasets reported in Section 3.1. In Table 1, we observe for both toddlers and adults that an image size of $128\times 128$ produces the best recognition accuracy for all gaze strategies. Importantly, Toddler and Adult gaze fixations $128\times 128$ present an accuracy boost of $8\%$ compared to Centroid gaze $480\times 480$, which simulates head-cameras recordings without an eye-tracker. We conclude that accounting for the constrained size of the central visual field is crucial for learning powerful object representations. We speculate that, this boost originates in the ability of a $128\times 128$ gaze-centered crop to capture the complete structure of an object while minimizing irrelevant background information.
@@ 4.2 Constrained size of the central visual field is crucial for learning object representations
Table 1: Linear object recognition accuracy of different cropping sizes. We have bolded the main results of gaze fixation, while the underlined results represent simulations that do not utilize actual gaze fixation and consider only the egocentric visual experience.
@@
4.3 Toddlers’ gaze behavior boosts very slowly changing representations
@@
Previous work suggests that semantic aspects of the visual experience vary more slowly for toddlers than for adults (Sheybani et al., 2023) and that extending the gap of time between two positive pairs can improve the quality of object representations if visual inputs are sufficiently stable over time (Aubret et al., 2022; Schneider et al., 2021). Thus, we investigate whether amplifying the temporal slowness of our representation intensifies the difference between toddlers’ and adults’ representations. To amplify temporal slowness, we increase the temporal gap $\Delta{T}$ between representations that are made similar. As shown in Figure 5A, $\Delta{T}$ ranges from $\frac{1}{30}$ to 3.0 seconds, increasing continuously by 0.5 seconds at each step. The models trained with the Toddler fixation dataset achieve the highest recognition accuracy when $\Delta{T}=1.5$s. Conversely, Figure 5B shows that, for models trained with the Adult fixation dataset, increasing the interval between positive pairs decreases the quality of object representations. The results are consistent for both human fixations and centroid fixations (“Fixation” and “Centroid”). We conclude that even younger toddlers are expected to show even slower behavior compared to adults. This leads to a better visual representation for toddlers.
@@
Figure 5: The impact of different $\Delta{T}$ on recognition accuracy for A- toddlers and B- adults according to different fixation datasets.
4.4 Toddlers’ advantage hinges on longer inspections of objects when holding them
@@
So far, the temporal structure of toddlers’ egocentric experience clearly supports object learning better than adults. However, the temporal properties responsible for this effect remain unclear. Here, we further analyze the visual statistics of central visual experiences. We focus on four metrics that characterize the temporal sequence of images: the average fixation duration before making a saccade, the average duration of object looking bouts, the average duration of object looking when the camera-wearer holds the object, and the cumulative duration of object looking in a recording. We explain how we detect saccade and compute the average fixation duration in Appendix A. For other metrics, we leverage manually labeled timestamps (by (Bambach et al., 2018)) about when toddlers and adults look at/hold an object. In the following, we label ”Object looking” when the gaze fixation points are located on an object while the camera-wearer is not holding the object. We successfully extracted the data from 28 out of 38 toddlers and conducted all subsequent experiments using these 28 toddlers. The remaining participants are excluded from this section due to the lack of data on fixation duration. Table 2 in Appendix C presents the details of these specific 28 toddlers.
@@
Figure 6: Correlation analysis between the linear recognition accuracy and A- the average fixation duration; B- the average duration for object looking; C- the average duration during object looking while holding the object D- the cumulative duration for object looking. Models were all trained on individual Toddler and Adult fixation datasets. In each figure, the crosshairs represent the mean and standard deviation of the data values over the two axes. The legends show the Pearson correlation coefficients and their p-values.
@@
Figure 7: Comparison of A- average fixation duration; B- average duration for object looking; C- average duration for object looking and holding between toddlers and adults. Each panel includes the frequency distribution for the given metric, along with a density curve. The t-test statistics and p-values are given in the titles.
@@
In Figure 6, we observe that object recognition accuracy is highly linearly correlated with the three average durations but only weakly correlated with the cumulative duration of object looking. This indicates that long fixation bouts are important in explaining the quality of visual representations trained on the Toddler and Adult fixation datasets.
@@ 5 Conclusion
Current SSL approaches still struggle to learn robust human-like object representations and the reasons for this remain unclear. Here, we investigated whether a biologically inspired visual learning model can take advantage of toddlers’ gaze behavior to develop robust object representations. We cropped the toddlers’ gaze location from egocentric and eye-tracking recordings during play sessions. Then, we trained a bio-inspired unsupervised model that drives visual representations to slowly change. Our findings indicate that toddlers’ gaze strategies boost the learning of representations that support view-invariant object instance recognition within a single play session of 12 minutes. In particular, the effect was weaker with adults’ gaze behavior. Our analysis shows that our approximated central visual experience is crucial for learning object-oriented representations and that toddlers’ gaze behavior makes better use of slowness-based learning compared to adults. The latter point is supported by looking longer at objects while holding them. During their relatively long holding periods, we speculate that toddlers may actually be turning and moving the object, giving access to high-quality sequences that contain different object views over a short period of time.
From a developmental perspective, our work provides strong evidence that the development of view-invariant representations can originate in slowly changing representations, a mechanism previously found in the brain (Li & DiCarlo, 2008; Miyashita, 1988). We further demonstrate that toddlers may curate their gaze behavior to enhance the quality of their visual representations. However, it remains unclear if toddlers intentionally optimize their learning progress in visual representations. From a machine learning perspective, we show that combining eye-tracking video data and SSL supports unsupervised view-invariant recognition. This work marks a significant step towards learning strong representations without hand-crafted image datasets (e.g., (Aubret et al., 2022)).
@@
Acknowledgments
ZY thanks the Xidian-FIAS International Joint Research Center for funding. Additional support was received by the Deutsche Forschungsgemeinschaft (DFG project 5368 “Abstract REpresentations in Neural Architectures (ARENA)”), as well as the project “The Adaptive Mind” funded by the Excellence Program of the Hessian Ministry of Higher Education, Science, Research and Art (HMWK). Jochen Triesch was supported by the Johanna Quandt foundation. We gratefully acknowledge the Goethe-University (NHR Center NHR@SW) for providing computing and data-processing resources needed for this work.
@@ Appendix A Additional details
In section 3.1, we explain that the center of the frames is misaligned with respect to the stationary position of the eyes. To support this statement, Figure 8 and Figure 9 display the distribution of gaze locations for each toddler and adult, respectively. Brighter areas indicate higher frequencies of gaze fixation at those locations. The results indicate their average gaze location is not centered with respect to the camera.
@@
Appendix B Complementary analysis
B.1 Robustness testing with varying hyper-parameters
@@
Figure 11: The impact of A. average saccade duration and B. toddler’s age on recognition accuracy.
B.2 Study of saccade duration and age
@@
Appendix C Details of all toddlers’ Data
We provide information for all toddlers participating in the study in Table 2. The toddler ID represents an anonymized identifier for each toddler.

### v2

@@ 4 Results
We aim to investigate whether toddlers’ gaze behavior during a play session supports learning view-invariant object recognition. We also want to analyze the factors contributing to this.
@@
4.1 Toddlers’ central visual field experience supports the learning of invariant object representations via time-based SSL
To test if a toddler’s gaze behavior supports the learning of strong object representations, we compare the representations learned by SimCLR-TT when trained on the different datasets introduced in Section 3.1 (Figure 4). Results for BYOL-TT (Schneider et al., 2021) show a similar trend and are given in Appendix B.1. We find that models trained with the Toddler fixation dataset outperform those trained with the Random fixation dataset (toddler) or the Centroid fixation dataset (toddler). This suggests that biologically inspired visual learning models like SimCLR-TT can leverage human gaze behavior to learn invariant object representations.
We wondered whether the visual experience of only a single human during a play session suffices to build good visual representations. To investigate this question, we train SimCLR-TT on the individual ecordings of each toddler and adult separately and compute the average of linear accuracies. We train the encoder (ResNet18) using all fixation data from a single toddler/adult, followed by training and testing the linear classifier with the Objects fixation data from the same and different toddlers/adults. We control the training set to comprise 75% of the total data, ensuring that the test set does not overlap with the training set. Figure 4 shows that the central visual experience of one toddler leads to representations almost as good as those from the central visual experiences of all toddlers. We show additional results with a larger ResNet50 in Appendix B.2.
Finally, we assess whether toddlers’ visual experience produces better or worse representations than that of adults. By comparing the object recognition accuracy of models trained on fixation datasets from toddlers and adults, we see the same results. Toddlers’ experiences induce more robust representations compared to adults when training with one toddler/adult trial, as well as when training with the entire dataset. We conclude that, toddlers’ central visual experience supports more data-efficient learning than adults. Overall, toddlers appear to successfully curate their gaze behavior to permit the learning of robust object representations.
Figure 4: Linear object recognition accuracy of training on individual participants and all participants. We use red and blue to represent the test results of the model trained on the relevant datasets for toddlers and adults, respectively. The two gray bars represent the Oracles and indicate the test results on the Plain background and Object fixation datasets. Specifically, in the experiment, Oracles refers to the Objects fixation dataset and the Plain background dataset. For gaze fixation, we compared the test results of the model trained on each participant and across all participants. Specifically, the mean accuracy is shown, when trained with 38 dyads individually. For the all-participants setting, 30 participants are randomly selected from the pool of 38 for the training set under each random seed. The vertical bars represent the standard deviation over three random seeds.
4.2 Constraining input to the central visual field improves learning
[…]
@@ 4.2 Constraining input to the central visual field improves learning
Table 1: Linear object recognition accuracy for different cropping sizes. We have bolded the main results of gaze fixation, while the underlined results represent simulations that do not utilize actual gaze fixation and consider only the egocentric visual experience.
@@
4.3 Toddlers’ gaze behavior favors stronger emphasis on slowness
@@
Previous work suggests that semantic aspects of the visual experience vary more slowly for toddlers than for adults (Sheybani et al., 2023) and that extending the gap of time between two positive pairs can improve the quality of object representations if visual inputs are sufficiently stable over time (Aubret et al., 2022; Schneider et al., 2021). Thus, we investigate whether amplifying the temporal slowness of our representation intensifies the difference between toddlers’ and adults’ representations. To amplify temporal slowness, we increase the temporal gap $\Delta{T}$ between representations that are made similar. As shown in Figure 5A, $\Delta{T}$ ranges from $\frac{1}{30}$ to 3.0 seconds, increasing continuously by 0.5 seconds at each step. The models trained with the Toddler fixation dataset achieve the highest recognition accuracy when $\Delta{T}=1.5$s. Conversely, Figure 5B shows that, for models trained with the Adult fixation dataset, increasing the interval between positive pairs decreases the quality of object representations. The results are consistent for both human fixations and centroid fixations (“Fixation” and “Centroid”). We conclude that toddlers’ gaze behavior favors a stronger emphasis on slowness (greater $\Delta T$) than that of adults.
@@
Figure 5: The impact of different $\Delta{T}$ on recognition accuracy for toddlers (A) and adults (B).
4.4 Toddlers’ long object inspections relative to adults facilitate learning
@@
So far, we have shown that the egocentric visual experience of toddlers facilitates the self-supervised learning of object representations relative to that of adults. However, the temporal properties responsible for this effect remain unclear. Here, we further analyze the visual statistics of central visual experiences. We focus on four metrics that characterize the temporal sequence of images: the average fixation duration before making a saccade, the average duration of object looking bouts, the average duration of object looking when the camera-wearer holds the object, and the cumulative duration of object looking in a recording. We explain how we detect saccades and compute average fixation durations in Appendix A. For other metrics, we leverage manually labeled timestamps (by (Bambach et al., 2018)) about when toddlers and adults look at/hold an object. In the following, we label “Object looking” when the gaze fixation points are located on an object while the camera-wearer is not holding the object. We successfully extracted the data from 28 out of 38 toddlers and conducted all subsequent experiments using these 28 toddlers. The remaining participants are excluded from this section due to the lack of data on fixation durations. Table 3 in Appendix C presents the details of these specific 28 toddlers.
@@
Figure 6: Correlation analysis between the linear recognition accuracy and the average fixation duration (A), the average duration of object looking (B), the average duration of object looking while holding the object (C), and the cumulative duration of object looking (D). Models were all trained on individual Toddler and Adult fixation datasets. In each figure, the crosshairs represent the mean and standard deviation of the data values over the two axes. The legends show the Pearson correlation coefficients and their p-values.
@@
Figure 7: Comparison of average fixation duration (A), average duration of object looking (B), average duration of object looking while holding (D) for toddlers and adults. Each panel includes the frequency distribution for the given metric, along with a density curve. The t-test statistics and p-values are given in the titles.
@@
In Figure 6, we observe that object recognition accuracy is highly correlated with the three average durations but only weakly correlated with the cumulative duration of object looking. This indicates that long fixation bouts are important in explaining the relative quality of visual representations trained on the Toddler vs. Adult fixation datasets.
@@ 5 Conclusion
Current SSL approaches still struggle to learn robust human-like object representations and the reasons for this remain unclear. Here, we investigated whether biologically inspired visual learning models can take advantage of toddlers’ gaze behavior to develop robust object representations. We cropped the toddlers’ gaze location from egocentric video recordings with eye-tracking during play sessions. Then, we trained bio-inspired unsupervised models that drive visual representations to slowly change. Our findings indicate that toddlers’ gaze strategies permit the learning of representations that support view-invariant object instance recognition within a single play session of 12 minutes. Results were weaker for adults’ gaze behavior. Our analysis shows that our approximated central visual experience is crucial for learning object-oriented representations and that toddlers’ gaze behavior favors a stronger emphasis on slowness compared to adults. This is consistent with toddlers looking longer at objects while holding them. During their relatively long holding periods, toddlers may turn and move the object, giving access to high-quality sequences containing different object views over a short period of time.
From a developmental perspective, our work provides strong evidence that the development of view-invariant representations can originate from a slowness learning objective, a mechanism supported by neuroscientifc studies (Li & DiCarlo, 2008; Miyashita, 1988). We further demonstrate that toddlers may curate their gaze behavior to enhance the quality of their visual representations. From a machine learning perspective, we show that combining eye-tracking video data and SSL supports unsupervised view-invariant recognition. This work marks a significant step towards learning strong representations without hand-crafted image datasets (e.g., (Aubret et al., 2022)).
@@ Appendix A Additional details
In section 3.1, we explain that the center of the frames is misaligned with respect to the stationary position of the eyes. To support this statement, Figure 8 and Figure 9 display the distribution of gaze locations for each toddler and adult, respectively. Brighter areas indicate higher frequencies of gaze fixation at those locations. The results indicate that their average gaze location is not centered with respect to the camera.
@@
Appendix B Complementary analysis
B.1 Results of training BYOL-TT
In order to evaluate whether our conclusions also hold for different methods learning with temporal slowness, we perform the same experiments described in Section 4.1 with BYOL-TT. Similar to SimCLR-TT, BYOL-TT was originally considered to be used for contrastive learning through time (Schneider et al., 2021). Its loss function is defined as
$\mathcal{L}_{\theta_{t},\xi_{t+\Delta{T}}}=2-2\cdot\operatorname{sim}\left(\mathrm{q}_{\theta_{t}}(\mathrm{z}_{\theta_{t}}\right),\mathrm{z}_{\xi_{t+\Delta{T}}}),$ (2)
where $q_{\theta_{t}}(z_{\theta_{t}})$ is the prediction of the online network for one frame, $z_{\xi_{t+\Delta{T}}}$ represents outputs from the target network. Here, $\theta$ corresponds to the weights of the online network, and $\xi$ represents the weights of the target network. Again, we use the cosine similarity as the similarity function.
In Figure 10A, we found, in line with (Schneider et al., 2021) that BYOL-TT, as the backbone model, extracts less effective representations from the different fixation strategy datasets compared to SimCLR-TT. However, the relative relationships between the data remain unchanged. Overall, the conclusion that toddler fixation contributes to the acquisition of more robust representations still holds.
Figure 10: Linear object recognition accuracy of different settings. A. Testing results of BYOL-TT training on different datasets. B. Different backbone training on Toddler and Adult fixation dataset.
Figure 11: Sampling on Objects fixation dataset. A. Undersampling; B. Oversampling; C. Comparison of dataset sizes after using different sampling methods.
B.2 Impact of changing the self-supervised learning encoder
We compared the accuracy of BYOL-TT and SimCLR-TT using ResNet-18 and ResNet-50 as encoders on both the Toddler and Adult fixation datasets. As shown in Figure 10B, introducing more complex encoders resulted in a significant improvement in accuracy, with the gap between toddler and adult performance narrowing. This suggests that a more sophisticated encoder can equalize different boosting sampling strategies, which may obscure the inherent differences in representations between toddlers and adults. In contrast, a simpler encoder tends to profit more from toddler gaze behavior compared to those from adults.
B.3 Analysis of the class imbalance in the Objects fixation dataset
To investigate the impact of the imbalance in the Objects fixation dataset shown in Figure 2, we adjusted the distribution of the Objects fixation dataset while keeping the original encoder training results unchanged. The linear classifier was then trained and tested on the adjusted datasets. The number of categories remained fixed at 24 throughout the experiments. We compared the results of two types of sampling strategies:
Undersampling. We applied random undersampling to reduce the number of samples in the top 5 categories, making their quantities similar to those of the other categories. We do not intend to equalize all classes. In real-world scenarios, toddlers naturally show preferences for certain toys, and this behavior should be preserved. Our goal is to smooth the occurrence probabilities of other objects relatively rather than enforce an artificial balance across all categories.
Oversampling. Similarly, we applied random oversampling to increase the number of samples in the underrepresented categories to match the quantity of the top 5 categories. However, this method will result in duplicate samples in the dataset.
The data distributions after applying both sampling methods are shown in Figure 11. We maintain the experimental setup consistent with Section 4.1 and train a linear classifier on the undersampling and oversampling object fixation datasets.
In Figure 12, we observe that when the total sample size is reduced, the recognition accuracy of the models trained on Toddler and Adult fixation datasets decreases, but the difference in their accuracy continues to widen. However, with more complex or balanced training, the model’s generalization capacity improves, and the performance across toddlers and adults tends to converge, reducing the impact of differences in visual behaviors. Therefore, toddler gaze behavior might offer a greater advantage under undersampling conditions.
Figure 12: Linear object recognition accuracy and the difference in accuracy between undersampling and oversampling. A. We compared the recognition accuracy under different sampling methods, where “///” represents undersampling and “\” represents oversampling. Additionally, we provide the test results after training on one dyad versus all dyads; B. The difference in recognition accuracy between toddlers and adults under different sampling methods. Here, we also compare the accuracy differences of the model trained on one dyad versus all dyads.
Figure 13: Some evidence highlights the differences between toddlers and adults. In (A-B), We observe variations in the test accuracy of models training on the Toddler and Adult fixation datasets under different $\Delta{T}$. We attached fitted regression lines, and the shaded areas show the 95% confidence interval. (C-E) illustrates box plots showing the data differences between toddlers and adults across three metrics. The red line indicates the median value (Q2), and the gray dots represent outlier data exceeding the upper quartile (Q3).
B.4 Highlights differences between toddlers and adults
We provide additional evidence highlighting the differences between toddlers and adult. In Figure 13A and Figure 13B, we compare the changes in recognition accuracy for both toddlers and adults under different $\Delta{T}$ values. From the regression lines, the increasing $\Delta{T}$ amplifies the difference in recognition accuracy between training on toddlers’ and adults’ fixation datasets, consistent with the findings in Section 4.3. Besides Figure 7, Figure 13C-E display the box plots for the three corresponding metrics, revealing significant distinctions in the way toddlers and adults observe objects across all three metrics.
@@ B.4 Highlights differences between toddlers and adults
Figure 15: The impact of (A) average saccade duration and (B) toddler’s age on recognition accuracy.
B.5 Robustness testing with varying hyper-parameters
@@
B.6 Study of saccade duration and age
@@
Appendix C Details of all toys and toddlers Data
We provide information for all toys and toddlers participating in the study in Table 2 and Table 3. The toddler ID represents an anonymized identifier for each toddler.
Table 2: 24 toys were used for toddler interaction. Among them, “ Library ” refers to those toys that were successfully recognized when the toddler calls any word from the corresponding row. However, these columns are not within the scope of the current study’s discussion. The main focus is on the colors, shapes, or textures of these 24 toys, which are more likely to help toddlers differentiate between them.
Gazetag Naming ICONS ID Library
helmet [Uncaptioned image] 1 helmet, hat
house [Uncaptioned image] 2 house, home
bluecar [Uncaptioned image] 3 car
rose [Uncaptioned image] 4 rose, flower, plant
elephant [Uncaptioned image] 5 elephant
snowman [Uncaptioned image] 6 snowman
rabbit [Uncaptioned image] 7 rabbit, bunny
spongebob [Uncaptioned image] 8 spongebob, block
turtle [Uncaptioned image] 9 turtle, tortoise
hammer [Uncaptioned image] 10 hammer, tool, mallet
ladybug [Uncaptioned image] 11 bug, insect, ladybug, beetle
mantis [Uncaptioned image] 12 bug, insect, praying mantis, mantis, grasshopper
greencar [Uncaptioned image] 13 car
saw [Uncaptioned image] 14 saw, tool
doll [Uncaptioned image] 15 baby, baby doll, girl, doll
phone [Uncaptioned image] 16 phone, telephone
rubiks [Uncaptioned image] 17 block, rubiks cube, rubiks, cube
shovel [Uncaptioned image] 18 rake, shovel, tool
bigwheels [Uncaptioned image] 19 truck, jeep, bigwheel, car
whitecar [Uncaptioned image] 20 car, policecar
ladybugstick [Uncaptioned image] 21 ladybug, bug
purpleblock [Uncaptioned image] 22 block, cube
bed [Uncaptioned image] 23 bed
clearblock [Uncaptioned image] 24 block, cube